OpenStudy (katielong):

Factorising: i don't understand the method to answer the following... simplify (3p^2-21p)/(3p^2-20p-7) (i will put the actual equations below)

OpenStudy (katielong):

\[\frac{ 3p ^{2}-21p }{ 3p ^{2} -20p-7}\] i know that the top of the fraction would be: \[3p(p-7)\] but what is the bottom?

Directrix (directrix):

The denominator factors as (3p+1)(p-7)

OpenStudy (katielong):

okay thanks, how do you find this out? that's just the part i'm struggling with :L

OpenStudy (katielong):

eg. how do you know it's (3p+1)(p-7) rather than (3p-7)(p+1)?

Directrix (directrix):

(3p-7)(p+1) does not multiply back to the correct middle term of -20p, -7 *p +1*(3p) = -7p + 3p = -4p and not -20p.

Directrix (directrix):

(3p-7)(p+1) does not equal (3p^2-20p-7)

Directrix (directrix):

(3p+1)(p-7) = (3p^2-20p-7)

OpenStudy (katielong):

ah okayy thanks, so another question: \[3x ^{2}+8x-3\]

Directrix (directrix):

Here's a technique for factoring trinomials I wrote up based on how I learned to factor by busting the "b." You might want to look over it later. Busting the "b" 4 x² -1x + 18 a = 4 and c = 18 and b = -1. The task on "bust the b" is to find numbers that mutiply to a*c AND add to b. In this case, numbers that multiply to 4*18 and add to -1. 14 * 18 breaks aparts into the product of prime factors : 2*2*2*3*3 Look at those five factor of 14 * 18 and try to put them in the form of 2 numbers that differ by -1. 2*2*2 and 3*3 appear to differ by 1. That would be 8 and 9. "b" has been "busted into 8 and - 9 whose sum is -1 which equals "b" in this problem. 4x - x+ 18 = 4 x² + 8 x - 9 x - 1. The "b" has been busted. Factoring by grouping comes next. 4 x² + 8 x - 9 x - 18 = 4x ( x + 2) - 9 ( x + 2 ) = --> On this step, (x + 2) is the common factor of the expression. [ (x + 2) ] ( 4x - 9 ) = ( x + 2 ) ( 4x - 9)

Directrix (directrix):

Did you finish this problem: simplify (3p^2-21p)/(3p^2-20p-7)

Directrix (directrix):

If so, what did you get for the answer?

Directrix (directrix):

Is \[3x ^{2}+8x-3\] = 3x^2 + 8x - 3. The math processor is not working on the site or on my computer.

OpenStudy (katielong):

\[ \frac{ 3p }{ 3p+1 }\]

Directrix (directrix):

Click on the attachment to see what I see when the math processor is not working. :)

OpenStudy (katielong):

ah i get you... well basically it's: 3p/(3p+1)

Directrix (directrix):

3p/(3p+1) is the answer I got.

Directrix (directrix):

What are we doing here, factoring? \[3x ^{2}+8x-3\]

Directrix (directrix):

Look for two numbers that multiply to -9 and add to -8.

OpenStudy (agent0smith):

@Directrix i was just about to explain that, how to factor trinomials when a>1 :P I'll just link to my explanation, but yours is already here... but i do it slightly differently. http://www.wyzant.com/answers/4707/explain_how_to_factor_the_following_trinomials_forms_x_bx_c

OpenStudy (katielong):

Thanks for your help :) i will read it all through, if theres any more questions i'll pop them on here, tah :)

Directrix (directrix):

While we are here, we can factor 3x^2 + 8x -3 if you have time.

OpenStudy (katielong):

ah yeh thanks, so to start off: would it be (3x ?)(x ?)

OpenStudy (agent0smith):

First find factors of -9 (since 3*(-3) = -9) that add up to +8.

Directrix (directrix):

3x^2 + 8x -3 = 3x^2 + 9x - 1x -3 = 3x(x + 3) -1 (x+3)= (x + 3)(3x-1). It is factoring by grouping although bust the b comes from writing the 8x as the sum of 9x and -x.

OpenStudy (katielong):

the busting the b method seems way to hard, im doing gcse's year 11

Directrix (directrix):

It's just a slightly more organized technique than guess and check but that doesn't mean that busting the b is better than any other method. Whatever works for the learner is best, I think.

OpenStudy (katielong):

okay, well i will try a couple more questions. how about if its not a trinomial, eg 9x^2-1

OpenStudy (agent0smith):

The way i do it is similar to "busting the b", but w/o the grouping. factors of -9 that add up to 8 are -1 and 9, so i write: (3x -1)(3x+9) .... but when we multiplied 3*(-3) earlier, we multiplied our expression by 3. So divide it by 3: (3x -1)(3x+9)/3 = (3x-1)(x+3) Done!

Directrix (directrix):

There are a few special factoring patterns that can be automated to help with the difference of two squares and perfect square trinomials.

OpenStudy (agent0smith):

oops noticed an error... fixed. @katielong it'll get much easier with practice. I've seen students who really struggle with algebra use this method well. eg. 4x^2 + 10x + 4 .... first multiply the first 4 by the last 4. 4x4 = 16. Factors of 16 that add up to the 10... 8 and 2. So write (4x+8)(4x+2) ... and now don't forget to divide it all by the 4 we originally multiplied by: (4x+8)(4x+2)/4 = (x+2)(4x+2)

OpenStudy (katielong):

question attempt: 3c^2-c-10 3*10 and +'s to make 1? 30=3*5*2? so 3 and 2 have a difference of 1? 3c^2-c-10= 3c^2+3c-2c+1? nope im stuck..... :S

OpenStudy (katielong):

@agent0smith i see now... i did it wrong too... let me try again...

OpenStudy (agent0smith):

Don't worry Katie, you'll get it with practice. Well the way i do it is to just find factors... 3x-10=-30. Factors of -30 that also add up to -1... +5 and -6 since 5-6=-1 (3c-6)(3c+5) now divide by the 3 we first multiplied by: (3c-6)(3c+5)/3 = (c-2)(3x+5)

Directrix (directrix):

9x^2-1 = ?

OpenStudy (katielong):

what about say 2n^2-8?

OpenStudy (agent0smith):

^there you can first pull a factor of 2, then use diff. of two squares. \[2n^2-8 = 2(n^2 - 4) \]

OpenStudy (katielong):

ah i see thanks

OpenStudy (katielong):

so right, simplify (n^2+6n-16)/(9x^2-1) the top is (n+8)(n-2) the bottom is 2(n-4) so how do you do (n+8)(n-2)/2(n-4)?

OpenStudy (agent0smith):

wait... the bottom is 9x^2-1 which isn't 2(n-4). \[ 9x^2-1 = (3x)^2-1^2 \]now try using diff. of two squares on that. See directrix's pic above. I need sleep, will help later if you need it.

OpenStudy (katielong):

yeah thats fine thanks, haha its mid-day here :L

OpenStudy (agent0smith):

3:30 am here :P btw that one doesn't appear to simplify easily.... (n+8)(n-2)/(9n^2-1) diff of two squares on the bottom will give: \[\frac{(n+8)(n-2) }{ (3n-1)(3n+1) }\] so there doesn't seem to be common factors...

OpenStudy (katielong):

hang on 2 questions are muddled here... theres 1) \[\frac{ n^2+6n-16 }{ 2n^2-8 }\] 2) \[\frac{ 3x^2+8x-3 }{ 9x^2-1 }\]

OpenStudy (agent0smith):

ohhh... :) yeah then the first one becomes (diff of two squares on denominator): \[\frac{(n+8)(n-2)}{2(n^2-4) } = \frac{(n+8)(n-2)}{2(n+2)(n-2) } \] now try simplifying that :) second (which we factored the top part of earlier) \[\frac{ 3x^2+8x-3 }{ 9x^2-1 } = \frac{ (3x-1 )(x+3) }{(3x-1)(3x+1) }\]

OpenStudy (katielong):

thanks i've got it now! go get some sleep haha, sorry fro keeping you up :L

OpenStudy (agent0smith):

You're welcome! And no prob :) gnight!